# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 1 Introduction
Based on these similarities, we hypothesize the existence of cascade phenomena in the collisionless dark matter flow (SG-CFD). However, due to the collisionless nature and long-range gravity, we also expect the cascade in SG-CFD to be unique and distinct from turbulence. This paper will develop and formulate the ideas of mass and energy cascades in SG-CFD with the help of large-scale cosmological simulations. Despite the path already explored by many pioneers, these ideas provide a new understanding of the two important aspects of gravitational collapse: the abundance and internal structure of haloes. The rest of the paper is organized as follows: Section 2 introduces the simulation and numerical data. Section 3 presents the cosmic energy evolution that demonstrates the decrease of the total energy of dark matter in an expanding background. Section 4 presents the mass and energy cascade in halo mass space that leads to the halo random walk, halo mass functions, and scaling laws in mass space. Section 5 presents the energy cascade in individual haloes that lead to the particle random walk, halo density profiles, and universal scaling laws in haloes, followed by the Section 6 discussing the potential constraints on dark matter particle mass.
2 N-body simulations and numerical data
The large-scale cosmological Illustris simulation (Illustris-1-Dark) (Nelson et al., 2015) was used to demonstrate and validate the concepts. Illustris is a suite of large-volume cosmological dark matter only and hydrodynamical simulations. The selected Illustris-1-Dark is dark matter only simulations that has a cosmological volume of a 106.5Mpc3 and 18203 DM particles for a high mass resolution. Each DM particle has a mass around $m_{p}=7.6\times 10^{6}M_{\odot}$. The gravitational softening length is around 1.4 kpc. The simulation has cosmological parameters of dark matter density $\Omega_{DM}=0.2726$, dark energy density $\Omega_{DE}=0.7274$ at $z=0$, and Hubble constant $h=0.704$. Some key parameters of N-body simulations are listed in Table 1.
For cross-validation, the cosmological N-body simulations carried out by the Virgo consortium were also used. A comprehensive description can be found in (Frenk et al., 2000; Jenkins et al., 1998). The SCDM simulation of the Virgo consortium focuses on matter-dominant dark matter-only simulations with a standard CDM power spectrum (SCDM). The same set of simulation data has been widely used in a number of different studies, from clustering statistics (Jenkins et al., 1998) to the formation of halo clusters in large-scale environments (Colberg et al., 1999), and testing models for halo abundance and mass functions (Sheth et al., 2001). This simulation has a lower mass resolution with particle mass $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$. The simulation box is around 240 Mpc/h, where h is the dimensionless Hubble constant in the unit of 100 km/s/Mpc.
The friends-of-friends algorithm (FOF) was used to identify all haloes in each simulation that depend only on a dimensionless parameter b, which defines the linking length $b\left({N/V}\right)^{{-1/3}}$, where $V$ is the volume of the simulation box. In this work, haloes were identified with a linking length parameter of $b=0.2$. By identifying all haloes of different sizes, all dark matter particles were divided into halo particles with a total mass $M_{h}$ and out-of-halo particles that do not belong to any halo. Therefore, $M_{h}$ is the total mass of all haloes. We focus on the evolution of mass and energy in haloes of different mass $m_{h}$. All haloes were grouped into halo groups of various sizes according to the halo mass $m_{h}$ (or $n_{p}$, the number of particles in the halo), where $m_{h}=n_{p}m_{p}$. The total mass for a halo group of mass $m_{h}$ is $m_{g}=m_{h}n_{h}$, where $n_{h}$ is the number of haloes in that group. Two different cosmological simulations were used to demonstrate the fundamental concepts. The same approach can be easily extended to other cosmological simulations.
Table 1: Virgo (SCDM) and Illustris simulation (Illustris-1-Dark) parameters
Run $\Omega_{0}$ $h$ L (Mpc/h) N $m_{p}$ $M_{\odot}/h$ $l_{soft}$ (Kpc/h)
SCDM 1.0 0.5 239.5 $256^{3}$ 2.27×$\times 10^{11}$ 36
Illustris 0.24 0.704 75 $1820^{3}$ 5.28×$\times 10^{6}$ 1.4
3 Cosmic energy evolution
To better understand the cascade phenomenon, we first provide the energy evolution in self-gravitating collisionless dark matter flow. The equations of motion for N collisionless particles in comoving coordinates $\boldsymbol{\mathrm{x}}$ and physical time t read (Peebles, 1980):
where N particles have equal mass $m_{p}$. The Hubble parameter $H\left(t\right)={\dot{a}/a}$. Here, $H$ has a "damping" effect, which leads to the decrease in total energy of the N-body system (Eq. (6) and Fig. 1). In the matter-dominant era, the Hubble parameter satisfies $Ht=2/3$.
Next, we will derive the energy evolution based on equations of motion (Eq. (1)). We first introduce a transformed time variable s as ${ds/dt}=a^{p}$, where p is an arbitrary exponent. In terms of the new time variable $s$, the original Eq. (1) can be transformed to
where $\boldsymbol{\mathrm{F}}_{i}$ is the resultant force on particle i from all other particles, while $P_{s}$ is the total specific potential energy in comoving coordinates. Equation (2) reduces to the original Eq. (1) when $p=0$. With $p=-2$, the first-order derivative vanishes in Eq. (2) and s is the time variable for integration in N-body simulations. By setting $p=-1$, s is the conformal time. By setting $p=-{3/2}$ along with $H_{0}^{2}=H^{2}a^{3}$ for the matter-dominant era, the equation of motion becomes
In this transformed equation, the scale factor a is not explicitly involved. The time-dependent Hubble parameter $H$ is replaced by a Hubble constant $H_{0}$. This transformation offers significant convenience in analytically solving the cosmic energy evolution (Xu, 2022a).
We first identify the transformation between velocity $\boldsymbol{\mathrm{v}}_{i}$ in time variable $s$ and the peculiar velocity $\boldsymbol{\mathrm{u}}_{i}$,
The energy evolution of the N-body system can be obtained by multiplying ${\boldsymbol{\mathrm{v}}_{i}=d\boldsymbol{\mathrm{x}}_{i}/ds}$ on both sides of Eq. (2) and adding the equation of motion for all particles together (Xu, 2022a). An exact and simple equation (in time variable $s$) for the specific kinetic energy $K_{s}$ and the potential energy $P_{s}$ can be obtained as
$\frac{dK_{s}}{ds}+2HK_{s}(p+2)a^{-p}+a^{-(3+2p)}\frac{dP_{s}}{ds}=0.$ (5)
By setting $p=0$ and using the relations in Eq. (4), the exact cosmic energy equation for energy evolution of the N-body system reads
$\frac{\partial E_{y}}{\partial t}+H\left(2K_{p}+P_{y}\right)=0,$ (6)
which describes the energy evolution in an expanding background. Here $K_{p}$ is the peculiar kinetic energy, $P_{y}$ is the potential energy in physical, coordinates and $E_{y}=K_{p}+P_{y}$ is the total specific energy. This is also known as the Layzer-Irvine equation (Irvine, 1961; Layzer, 1963).
The cosmic energy evolution (Eq. (6)) admits a linear solution
where $\varepsilon_{u}<0$ is an important physical constant (unit: m2/s3) and the focus of this paper. It represents the rate of time variation of cosmic energy. The solution in Eq. (7) can be directly validated by N-body simulations. Figure 1 presents the energy evolution from Virgo simulation (SCDM). The (specific) kinetic energy $K_{p}$ and potential energy $P_{y}$ were calculated as the mean energy of all dark matter particles in the N-body system. Figure 1 confirms the solution in Eq. (7), i.e. a linear increase of $K_{p}$ with time and a negative potential energy $P_{y}=-1.4K_{p}$. Two key messages can be obtained:
i) According to the virial theorem, in virial equilibrium, $2K_{p}=n_{e}P_{y}$, where $n_{e}$ is an exponent of the interaction potential $r^{n_{e}}$. For standard gravity, exponent $n_{e}=-1$. Since gravity is the only interaction in the N-body system, we would expect $n_{e}=-1$ for an N-body system. However, the virial equilibrium can never be reached by the non-equilibrium N-body system (e.g., SG-CFD of dark matter flow). Instead of the virial equilibrium, a statistically steady state is established to continuously release system energy and maximize system entropy, which is manifested by the mass and energy cascade. In that particular state, the solution of Eq. (7) is valid such that an "effective" exponent $n_{e}=-10/7$ is obtained, which deviates from -1;
ii) The total energy $E_{y}=-0.4K_{p}$ decreases with time. In analog to hydrodynamic turbulence, the total energy of the N-body system in an expanding background decreases with time, as if "dissipated," even though there is no viscous force in the collisionless dark matter. What is the more fundamental nature of this "dissipation"? Will this energy "dissipation" lead to any dark radiation? These questions require further exploration (Xu, 2022b). Here, we admit that cosmic energy decreases with time and estimate the rate of energy "dissipation":
where $u_{0}\equiv u\left(t=t_{0}\right)\approx 350$km/s is the one-dimensional velocity dispersion of all dark matter particles, and $t_{0}\approx$13.7 billion years is the physical time at present epoch or the age of the universe. Different simulations may have slightly different values of $u_{0}$ due to different cosmological parameters. However, the rate of energy "dissipation" $\varepsilon_{u}\sim-10^{-7}m^{2}/s^{3}$ should be a good estimate.
Hereafter, the word "dissipation" stands for the decrease in total energy caused by the expanding background instead of the dissipation caused by the viscous force in turbulence. In the statistically steady state, the energy "dissipated" on small scales should balance the energy cascaded from large scales. Therefore, Equation (7) also suggests an inverse cascade of kinetic energy from small to large scales at a rate of $\varepsilon_{u}$ and a direct cascade of potential energy at a rate of $-1.4\varepsilon_{u}$. The total energy is directly cascaded to small scales at a rate of $-0.4\varepsilon_{u}$ to provide the energy that is "dissipated" on small scales due to the Hubble expansion. In the next sections, we will illustrate how the mass and energy cascade initiate and proceed and their impacts on structure formation and evolution.
4 Mass and energy cascade in halo mass space
4.1 Basic concepts
Figure 2: Schematic plot for the mass and energy cascade in halo mass space due to hierarchical structure formation of dark matter haloes. Haloes of mass $M$ merging with a single merger (free particles of mass $m$) leads to a mass flux to a larger scale $M+m$, that is, the halo of mass $M$ moving into the next mass sale $M+m$ after merging. For a given halo of mass $m_{h}$, a merge occurs with an average waiting time $\tau_{g}(m_{h},z)$. For a group of $n_{h}$ haloes of the same mass $m_{h}$, merge occurs with an average waiting time $\tau_{h}(m_{h},z)=\tau_{g}/n_{h}$ (Eq. (9)). Because haloes have finite mass and kinetic energy, continuous merging facilitates an inverse mass and kinetic energy cascade from small to large scales across haloes of different masses. The scale-independent mass and energy flux ($\varepsilon_{m}$ and $\varepsilon_{u}$ are independent of mass scale $m_{h}$) are expected in the propagation range ($m=m_{p}<m_{h}<m_{h}^{*}$). The mass and energy cascade rates become scale-dependent in the deposition range ($m_{h}>m_{h}^{*}$) (see Figs. 3, 4 for $\varepsilon_{m}$ and Fig. 8 for $\varepsilon_{u}$).
For collisionless dark matter, long-range gravity requires the formation of a broad spectrum of haloes to maximize the entropy of the self-gravitating collisionless system (Xu, 2023a). Highly localized halo structures are a major manifestation of non-linear gravitational collapse (Neyman and Scott, 1952; Cooray and Sheth, 2002). As the building blocks of dark matter flow (in contrast to the "eddies" in turbulence), haloes facilitate the mass and energy cascade across different halo masses. Figure 2 provides a schematic description of the mass and energy cascade in the halo mass space. This description is fully consistent with the hierarchical structure formation, where haloes grow via a series of merging events.
In Fig. 2, the merging of a halo (mass $M$) with a single merger of mass $m$ (depending on the mass resolution) leads to a mass and energy flux to a larger mass scale $M+m$, that is, the halo of mass $M$ moving into the next mass sale $M+m$ after merging. For a given halo of mass $m_{h}$, the merge event occurs with an average waiting time $\tau_{g}(m_{h},z)$. On average, haloes of mass $m_{h}$ need to wait for a specific waiting time $\tau_{g}$ before merging with a single merger. That waiting time is dependent on the mass of haloes. Larger haloes tend to accrete mass faster with a shorter waiting time ($\tau_{g}\propto m_{h}^{-\lambda}$ in Eq. (19)) with halo geometry parameter $\lambda=2/3$.
For a group of $n_{h}$ haloes of the same mass $m_{h}$, the average waiting time for a merge event of any halo in that group can be found as $\tau_{h}(m_{h},z)=\tau_{g}/n_{h}$. A simple explanation is provided here. Let us randomly pick one halo in that group to merge in every round and repeat until the given halo is picked. The probability of picking that given halo in the first round is $1/n_{h}$, and the corresponding waiting time for that specific halo is $\tau_{h}$. The probability of picking that given halo in the second round is $(n_{h}-1)/n_{h}*1/n_{h}$ (product of probability not selected in the first round and the probability selected in the second round). The corresponding waiting time is $2\tau_{h}$. Therefore, the waiting time $\tau_{g}$ for a given halo is
$\begin{split}\tau_{g}=\sum_{k=1}^{\infty}\frac{k\tau_{h}}{n_{h}}\left(\frac{n_{h}-1}{n_{h}}\right)^{k-1}=\frac{\tau_{h}}{n_{h}}+\frac{n_{h}-1}{n_{h}}\frac{\left(2\tau_{h}\right)}{n_{h}}+...=n_{h}\tau_{h}.\end{split}$ (9)
Based on this picture, the rate of mass cascade (or mass flux) in halo mass space can be conveniently defined as
$\begin{split}&\Pi_{m}(m_{h},z)=-\frac{m_{h}}{\tau_{h}}=-\frac{n_{h}m_{h}}{\tau_{g}}=-\frac{m_{g}}{\tau_{g}},\\ &\varepsilon_{m}(z)=\Pi_{m}(m_{h},z)\quad\textrm{for}\quad m_{h}\leq m_{h}^{*},\end{split}$ (10)
i.e., the mass of halo $m_{h}$ is passed to a larger mass scale during a period of $\tau_{h}$. Here, $m_{g}=n_{h}m_{h}$ is the total mass for $n_{h}$ haloes of the same mass $m_{h}$ (or the mass of the halo group) such that $\tau_{g}$ is the time required to pass the mass of the entire group $m_{g}$ to a larger mass scale. The negative sign represents the inverse cascade from small to large scales, in contrast to the direct cascade from large to small scales. The rate of the mass cascade is scale independent; that is, $\varepsilon_{m}$ is independent of the mass scale $m_{h}$ in the mass range $m_{h}<m_{h}^{*}$. The formulation of $\varepsilon_{m}$ in the mass cascade and the results of the N-body simulations are presented in Section 4.2 (Figs. 3 and 4).
Accompanied by the mass cascade, there exists also a simultaneous energy flux across haloes of different scales. To understand this, we will first demonstrate the energy "dissipation" that is intrinsically associated with every merging of haloes. This provides a small-scale mechanism for the "dissipation" of the total cosmic energy (Fig. 1). As energy continuously "dissipates" via halo merging, an energy cascade in halo mass space is required to balance the "dissipation". Here, we consider the two-body merging for haloes of mass $M_{1}$ and $M_{2}$, and size $r_{h1}$ and $r_{h2}$, respectively. After the merging, the new halo has a mass of $M_{3}=M_{1}+M_{2}$ and size $r_{h3}$. We calculate the difference in the total energy before and after the merging. From the virial theorem, the halo kinetic energy and potential energy are related as KE=$n_{e}$/2 PE. Therefore, the change in the total energy upon the merging reads
$\begin{split}\Delta E=-\left(1+\frac{n_{e}}{2}\right)\frac{GM_{3}^{2}}{r_{h3}}+\left(1+\frac{n_{e}}{2}\right)\left[\frac{G{M_{1}}^{2}}{r_{h1}}+\frac{G{M_{2}}^{2}}{r_{h2}}\right].\end{split}$ (11)
The density of haloes is related to the mean density $\bar{\rho}_{DM}$ as
@@ 4.1 Basic concepts
where $\Delta_{c}=18\pi^{2}$ is the critical density ratio from the spherical collapse model. Substituting Eq. (12) into Eq. (11), the change in the total energy due to the halo merging reads
$\small\begin{split}&\Delta E=\left(1+\frac{n_{e}}{2}\right)\left(\frac{4}{3}\pi\Delta_{c}\bar{\rho}_{DM}\right)^{\frac{1}{3}}G{M_{1}}^{\frac{5}{3}}\left(1+\gamma^{\frac{5}{3}}-(1+\gamma)^{\frac{5}{3}}\right),\end{split}$ (13)
where $\gamma=M_{2}/M_{1}$ is the mass ratio between two haloes. Clearly, the change in total energy due to the merging is always negative or $\Delta E<0$. Since a closed system tends to minimize its energy and maximize its entropy, the structure formation must proceed in a "bottom-up" fashion via halo merging to release system energy, not the "top-down" fashion. The halo merging provides a viable mechanism to "dissipate" the total cosmic energy in an expanding background. For two haloes with a large mass ratio or $\gamma=M_{2}/M_{1}\ll 1$,
$\begin{split}&\Delta E=-\frac{5}{3}\left(1+\frac{n_{e}}{2}\right)G\left(\frac{4}{3}\pi\Delta_{c}\bar{\rho}_{DM}\right)^{\frac{1}{3}}{M_{1}}^{\frac{2}{3}}M_{2}<0,\end{split}$ (14)
where the energy change is proportional to $M_{1}^{2/3}M_{2}$.
In halo mass space, the kinetic energy is cascaded from small to large mass scales at a rate of $\varepsilon_{u}$ (the same direction as the mass cascade). In contrast, the potential energy is cascaded in the opposite direction at a rate of $-1.4\varepsilon_{u}$ (Fig. 1). Therefore, the total energy is cascaded to small scales at a rate of $-0.4\varepsilon_{u}$. Similarly to the mass flux in Eq. (10), the flux of the specific kinetic energy across haloes reads (also see Eq. (40))
where $\overline{K_{pv}}$ is the mean specific kinetic energy in all haloes greater than $m_{h}$ (defined in Eq. (38)), while $\Lambda_{m}$ is the total mass in all haloes greater than $m_{h}$ (defined in Eq. (17)). For every waiting time of $\tau_{h}$, the energy cascade leads to an increase in the specific kinetic energy of $m_{h}{\overline{K_{pv}}}/{\Lambda_{m}}$ on all scales larger than $m_{h}$. The formulation of $\varepsilon_{u}$ in the energy cascade and the results of the N-body simulation are presented in Section 4.5 (Fig. 8).
In summary, the rate of the mass cascade $\varepsilon_{m}$, the rate of the energy cascade $\varepsilon_{u}$, and the waiting time $\tau_{g}$ are three central quantities that we will focus on. The mass and energy cascade establishes a statistically steady state to continuously release the energy of the system and maximize the entropy (Xu, 2023a). When the dark matter flow reaches that statistically steady state, the rates of mass and energy cascade must be scale independent (i.e., $\varepsilon_{m}$ and $\varepsilon_{u}$ are independent of mass scale $m_{h}$). If this is not the case, there would be a net accumulation of mass and energy on some intermediate-mass scale below $m_{h}^{*}$. The mathematical formulations of the mass and energy cascade are presented in Sections 4.2 and 4.5.
4.2 Formulating the mass cascade in mass space
In this section, we present a quantitative description of the mass cascade in the halo mass space. The energy cascade in the mass space can be formulated in a similar way (Section 4.5). First, we introduce the mass flux ($\Pi_{m}$) across haloes of different sizes
@@ 4.2 Formulating the mass cascade in mass space
The total mass in all haloes can be obtained by setting $m_{h}\rightarrow 0$ in Eq. (17), i.e. $M_{h}(t)=\Lambda_{m}(m_{h}\rightarrow 0,t)$. The time derivative of $\Lambda_{m}$ describes the mass flux from all haloes below the scale $m_{h}$ to all haloes above the scale $m_{h}$, i.e., the rate of mass cascade $\Pi_{m}$ in Eq. (16). In N-body simulations, we use the difference of $\Lambda_{m}$ at two different redshifts $z_{1}$ (or $t_{1}$) and $z_{2}$ (or $t_{2}$) to compute the time derivative and hence the mass flux $\Pi_{m}$.

### v4

@@ 1 Introduction
Based on these similarities, we hypothesize the existence of a universal range and associated "cascade" physics to facilitate the energy "dissipation" in the collisionless dark matter flow (SG-CFD). However, because of the collisionless nature and long-range gravity, we also expect the cascade in the SG-CFD to be unique and distinct from turbulence. Previous work has already studied universal scaling laws, the halo mass function, and density profiles (Xu:2023-Dark-matter-halo-mass-functions-and; Xu:2023-Universal-scaling-laws-and-density-slope). This paper will focus on the origin of universality, further developing and formulating these fundamental ideas with the help of cosmological simulations. Despite the path already explored by many pioneers, these ideas provide a new understanding of the abundance and internal structure of haloes.
The remainder of the paper is organized as follows: Section 2 introduces and formulates the universal transition range. Section 3 presents the mass and energy cascade in the halo mass space that leads to the halo random walk, halo mass functions, and universal scaling laws in the mass space. Section 4 presents the energy cascade in individual haloes and the particle random walk that leads to halo density profiles and universal scaling laws in haloes, followed by Section 5 discussing the constraints on dark matter particle mass. The appendix presents the evolution of the cosmic energy that is globally "dissipated" to enable the cascade in SG-CFD, and provides more details on the mass and energy cascade.
2 Universal transition spectrum with n=-1
The universal range between two critical scales $k_{NL1}$ and $k_{NL2}$ in Fig. 1 represents a transition between the linear and nonlinear regimes, with a universal spectrum index and an expanding range or ratio $k_{NL2}/k_{NL1}$. To find that spectrum index, we write a self-similar dimensionless spectrum $\Delta^{2}_{X}(k,a)$ around the characteristic scale $k_{h}^{*}$, which is inside that range. Here $k_{h}^{*}\sim 1/r_{hc}^{*}$ is the comoving scale associated with haloes of characteristic mass $m_{h}^{*}$. We can separate the time and k-dependence for the spectrum around $k_{h}^{*}$ as
$\Delta^{2}_{X}(k,a)=\frac{P(k,a)k^{3}}{2\pi^{2}}=\Delta^{2}_{X}(k^{*}_{h},a)\left(kr^{*}_{hc}\right)^{c}.$ (1)
where $c$ is the index to be determined. Here, the comoving halo size $r_{hc}^{*}$ is related to the comoving density of the halo $\rho_{hc}^{*}$ as
$\rho_{hc}^{*}=\frac{m_{h}^{*}}{\frac{4}{3}\pi{r_{hc}^{*}}^{3}}=\delta_{c}\bar{\rho}_{0}\quad\textrm{and}\quad H_{0}^{2}=\frac{8}{3}\pi G\bar{\rho}_{0},$ (2)
@@ 2 Universal transition spectrum with n=-1
where $\bar{\rho}_{0}$ is the mean density of matter at zz=0. The density ratio $\delta_{c}=18\pi^{2}$ can be obtained from the spherical collapse model for the matter-dominant era. Substituting (2) into (1) leads to
$\Delta^{2}_{X}(k,a)=\Delta^{2}_{X}(k^{*}_{h},a)\left(\frac{2Gm_{h}^{*}}{\delta_{c}H_{0}^{2}}\right)^{c/3}k^{c}.$ (3)
The time dependence of the spectrum comes from both $\Delta^{2}_{X}(k^{*}_{h},a)$ and the characteristic mass $m_{h}^{*}(a)$.
Figure 2: The power spectrum $P(k,z)$ at zz=0 from different Virgo simulations. Symbols plot the power spectrum with various cosmologies, exhibiting a universal transition range $P(k)\propto k^{-1}$ on scales $k_{NL1}<k<k_{NL2}$, where the gravitational dynamics dominate over the effect of the initial spectrum. Solid lines present the power spectrum from a power-law initial spectrum $P_{ini}(k)\propto k^{n_{i}}$, with $n_{i}$= -2, -1.5, -1, and 0. In the fully nonlinear range on scales $k>k_{NL2}$, the spectrum $P(k)$ retains the memory of the initial spectrum and develops an asymptotic slope $n_{a}$ as a function of the initial spectrum index $n_{i}$ (See Fig. 3).
When the universal spectrum is fully developed, we can relate the dimensionless spectrum to the two-point correlation function $\xi(r_{hc}^{*},a)$ on the same scale. Since nonlinear structures are always formed on the scale $r_{hc}^{*}$, the correlation function can be further related to linear overdensity $\delta_{vir}^{lin}\approx 1.69$. Therefore, we write
$\Delta^{2}_{X}(k_{h}^{*},a)\sim\xi(r_{hc}^{*},a)\sim(\delta_{vir}^{lin})^{2}a^{2}.$ (4)
In the spherical collapse model, the critical density ratio $\delta_{c}$ is related to the linear density $\delta_{vir}^{lin}$ as (Gunn:1972-Infall-of-Matter-into-Clusters; Xu:2022-Postulating-dark-matter-partic)
$\delta_{c}=32\xi_{ta}\quad\textrm{and}\quad\delta_{vir}^{lin}=\frac{3}{5}2^{2/3}(\xi_{ta})^{1/3},$ (5)
where $\xi_{ta}$ is the density ratio at the turn-around moment.Therefore, the linear overdensity and the dimensionless spectrum at $k_{h}^{*}$ read
$\delta_{vir}^{lin}=\frac{3}{10}\delta_{c}^{1/3}\quad\textrm{and}\quad\Delta^{2}_{X}(k_{h}^{*},a)\sim\delta_{c}^{2/3}a^{2}.$ (6)
Substituting into Eq. (3) leads to the fully developed spectrum
$\Delta^{2}_{X}(k,a)\propto(\delta_{c})^{\frac{2-c}{3}}a^{2}\left(\frac{Gm_{h}^{*}}{H_{0}^{2}}\right)^{c/3}k^{c}.$ (7)
Since gravitational dynamics and expansion are the only dominant physics in the transition region around scale $m_{h}^{*}$, relevant constants for the universal spectrum should include and only include the gravitational constant $G$, the characteristic mass $m_{h}^{*}$, and the Hubble parameter $H$. This requires cc=2. The spectrum is independent of $\delta_{c}$ and $a^{2}$ is absorbed into $H$ so that
$\Delta^{2}_{X}(k,a)\propto\left({Gm_{h}^{*}}/{H^{2}}\right)^{2/3}k^{2}.$ (8)
Note that this argument is independent of the cosmological parameters or the initial spectrum, so that the spectrum $P(k)\propto\Delta^{2}(k)/k^{3}\propto k^{-1}$ should be universal. In addition, the mass cascade theory also supports a $k^{-1}$ spectrum, as we will discuss in Section 3. The final expression can be conveniently written as (from Eq. (3))
$\Delta^{2}_{X}(k,a)=\Delta^{2}_{X}(k_{h}^{*},a)\left(\frac{2Gm_{h}^{*}H_{0}}{\delta_{c}(100\textrm{km/s})^{3}}\right)^{2/3}\left(k\frac{Mpc}{h}\right)^{2}.$ (9)
Figure 3: The variation of asymptotic slope $n_{a}$ of nonlinear spectrum $P(k)$ with the slope $n_{i}$ of initial power-law spectrum. Symbols plot the simulation results. The dashed line plots Peebles’ prediction without involving a universe transition range ($k_{NL1}=k_{NL2}$ and $\alpha_{x}=0$). The discrepancy with simulation data suggests a missing piece in that model. We propose that the existence of a universal range also impacts the asymptotic slope $n_{a}$, leading to a better agreement with simulations (Eq. (14)). Solid lines plots the prediction with a ratio $k_{NL2}/k_{NL1}=a^{\alpha_{x}}$, where $\alpha_{x}=1/2$ is preferred.
Figure 2 presents the power spectrum $P(k)$ at $z=0$ from different Virgo simulations (Frenk:2000-Public-Release-of-N-body-simul). Symbols plot the spectrum from various cosmologies, including the "standard" CDM with $\Omega_{m}=1$, $h=0.5$, and $\sigma_{8}=0.61$ (SCDM); "Lambda" CDM with $\Omega_{m}=0.3$, $\Lambda_{0}=0.7$, $h=0.7$, and $\sigma_{8}=0.9$ (LCDM); "Open" CDM with $\Omega_{m}=0.3$, $h=0.21$, and $\sigma_{8}=0.85$ (OCDM); and "Tau" CDM with $\Omega_{m}=1$, $h=0.5$, and $\sigma_{8}=0.61$ (tCDM). For clarity, the data were shifted (lowered) by one order of magnitude. A transition range can be clearly identified with a universal spectrum index of -1 that is independent of the cosmology model.
The effects of the initial power spectrum can best be studied by systematically varying a power-law initial spectrum. A more negative index $n$ often leads to less frequent halo merging and longer memory of the initial spectrum. Figure 2 also plots the simulated full spectrum with a power-law initial spectrum $P_{i}(k)\propto k^{n_{i}}$ (solid lines), with index $n_{i}$= -2, -1.5, -1 and 0, respectively (Smith:2003-Stable-clustering--the-halo-mo). The resultant spectrum $P(k)$ in the nonlinear range retains the memory of the initial spectrum and develops an asymptotic slope $n_{a}$ as a function of the initial index $n_{i}$.
Figure 3 presents the asymptotic slope $n_{a}$ as a function of the initial index $n_{i}$. Symbols plot the simulation data from Smith et al. (Smith:2003-Stable-clustering--the-halo-mo) and Widrow et al. (Widrow:2009-Power-spectrum-for-the-small-scale-Universe). Under the "stable clustering" hypothesis, Peebles shows that if the initial power spectrum was a pure power-law with an index $n_{i}$ in a matter dominant universe, the asymptotic slope $n_{a}$ of the nonlinear spectrum would be directly related to the initial index through the relation $n_{a}=-6/(5+n_{i})$ (Smith:2003-Stable-clustering--the-halo-mo) (dashed black line). However, the discrepancy between Peebles’ prediction and simulation data suggests that an expanding universal range is required.
We first divide the entire spectrum $\Delta^{2}(k)$ into linear range, universal transition, and nonlinear range.
$\begin{split}&\Delta^{2}_{L}(k,a)\propto a^{2}k^{3+n_{i}}\quad\textrm{for}\quad k\leq k_{NL1},\\ &\Delta^{2}_{X}(k,a)\propto a^{b}k^{2}\quad\textrm{for}\quad k_{NL1}\leq k\leq k_{NL2},\\ &\Delta^{2}_{NL}(k,a)\propto a^{-n_{a}}k^{3+n_{a}}\quad\textrm{for}\quad k\geq k_{NL2},\end{split}$ (10)
where the spectrum $\Delta^{2}_{NL}$ in the nonlinear range satisfies the stable clustering hypothesis (Smith:2003-Stable-clustering--the-halo-mo). The spectrum $\Delta^{2}_{NL}\propto a^{3}$ in the nonlinear range is due to fully bound structures on the physical scale with the average density of the universe $\propto a^{-3}$. For the universal range, $b$ is a parameter for time evolution that depends on the initial conditions, while the spectrum $\Delta^{2}_{X}\propto k^{2}$ is universal and independent of the cosmology and the initial spectrum. Since $k_{NL1}$ is determined by $\Delta^{2}_{L}(k_{NL1},a)\equiv 1$, we have the scale $k_{NL1}\propto a^{-2/(3+n_{i})}$.
The spectrum should be continuous at two critical scales $k_{NL1}$ and $k_{NL2}$, leading to
@@
Solving this continuity condition leads to the relation
@@
Next, we introduce a ratio parameter $\alpha_{x}>0$ to represent an expansion of the universal range